Shift Summary for evening shift, 22/06/99

Jianbo Wang (jianbo@fnal.gov)
Wed, 23 Jun 1999 00:03:48 -0500 (CDT)

Shift SummShift Summary for evening shift, 22/06/99

On Shift: Sasha G, Jianbo

Hours of beam delivered at all intensities: 7.3h
Hours of beam delivered at nominal intensity: 7.3h
Hours of E832 data (special runs not included): 7.2h
Hours of special runs (laser scans, muon runs...): muon 10min

Accounting of lost time:
A magnet quenched. It took ~ 20min to repair it.

========================================================================
Runs ended this shift:
Run 13692 - 1.1M events - BEAM_832
Run 13693 - 8.8M - BEAM_832
Run 13695 0.5M MUON
Run 13696 1.5M BEAM_832

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:

No

========================================================================
Overview of shift:
E832 datataking + muon run

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13692 15:40 BOA goes to 2.01, call MCR. Etot CsI in ktrig
window is red, reload ET map, it's still red.
SG reload ktrig window, the Etot CsI button is
green.

15:40 ktev2 is dead. reboot it.

16:02 Stop run 13692

13693 16:15 Start run 13693 Beam_832
Observed bad BOA (>2.3). Called the main control
room -- right at that moment magnet quenched/beam
lost

16:22 Etot is red in run init window, reload ET map.
Ktrig.tcl

16:30 Beam is back, BOA is 1.9. But it will be up and
down (1.6-2.3) for the whole run.

18:10 Ktevana errors: unknown crate in stream 5. In raw
data, the crate address B5 is missing. It should
be "FFFF XXB5", but it is "FFFF 0". B5 is trigger
latch crater. It happened three times in this
shift. Run Kfend, the problem is fixed.

19:00 Trigger scalar is out of limit, run kfrend, th
problem is gone. (Crate 33 DYC full)

19:15 One bad evt is found in e832 KQxYTF check
Eff=0.9996

147
148 19:45 Hcc inefficiency, channel 248.

19:46 SG:all these spills got bad BOA>2.0. Call MCR,
they said they were working on it.

20:05 ktev2 split tape window is red, mount QKNL03,
click, red button on tapewindow, the button turned
yellow, click on success, all butons turned green.

21:35 The tapes are full, run 13693 stopped.

13695 22:15 Start muon_4E12 run 13694
22:25 Stop run 13694

13696 22:25 Start Beam_832 run 13695. Polarity is negative
now.
22:10 Stop to update DC t0 (since expert is here: usual
mechanism of the beam time vasting).
13967 22:20 Start new run

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift